# Be More Diverse than the Most Diverse: Online Selection of Diverse Mixtures of Generative Models

### v1

@@ 4 Optimal Mixtures of Generative Models
Suppose we are given $m$ generative models, where model $i$ generates samples from the distribution $P_{i}$. If our goal is merely to find the model that minimize the loss (4), we should select $\mathrm{argmin}_{i}L(P_{i})$. Nevertheless, for diversity metrics such as RKE, it is possible that a mixture of the models will give a better diversity. Assume that the mixture weight of model $i$ is $\alpha_{i}\in[0,1]$, where $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{m})$ is a probability vector. The loss of the mixture distribution $\sum_{i=1}^{m}\alpha_{i}P_{i}$ can then be expressed as
$L(\boldsymbol{\alpha})$ $:=L\Big{(}\sum_{i=1}^{m}\alpha_{i}P_{i}\Big{)}=\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}+\mathbf{f}^{\top}\boldsymbol{\alpha},$
@@
Given $\mathbf{K},\mathbf{f}$, the probability vector $\boldsymbol{\alpha}$ minimizing $L(\boldsymbol{\alpha})$ can be found via a convex quadratic program.
In practice, we do not know the precise $\mathbf{K},\mathbf{f}$, and have to estimate them using samples. Suppose we have the samples $x_{i,1},\ldots,x_{i,n_{i}}$ from the distribution $P_{i}$ for $i=1,\ldots,m$, where $n_{i}$ is the number of observed samples from model $i$. Write $\mathbf{x}:=(x_{i,a})_{i\in[m],\,a\in[n_{i}]}$. We approximate the true mixture distribution $\sum_{i=1}^{m}\alpha_{i}P_{i}$ by the empirical mixture distribution $\sum_{i=1}^{m}\frac{\alpha_{i}}{n_{i}}\sum_{a=1}^{n_{i}}\delta_{x_{i,a}}$, where we assign a weight $\alpha_{i}/n_{i}$ to samples $x_{i,a}$ from model $i$, and $\delta_{x_{i,a}}$ denotes the degenerate distribution at $x_{i,a}$. We then approximate $L(\boldsymbol{\alpha})$ by the sample loss
$\hat{L}(\boldsymbol{\alpha};\mathbf{x})$ $:=L\Big{(}\sum_{i=1}^{m}\frac{\alpha_{i}}{n_{i}}\sum_{a=1}^{n_{i}}\delta_{x_{i,a}}\Big{)}=\boldsymbol{\alpha}^{\top}\hat{\mathbf{K}}(\mathbf{x})\boldsymbol{\alpha}+\hat{\mathbf{f}}(\mathbf{x})^{\top}\boldsymbol{\alpha},$ (5)
@@
The minimization of $\hat{L}(\boldsymbol{\alpha};\mathbf{x})$ over probability vectors $\boldsymbol{\alpha}$ is still a convex quadratic program.
@@ 5 Online Selection of Optimal Mixtures – Mixture Multi-Armed Bandit
If we are told by an oracle the optimal mixture $\boldsymbol{\alpha}^{*}$ that minimizes the loss $L(\boldsymbol{\alpha})$, then we should generate samples according to this mixture distribution, giving $\approx\alpha^{*}_{i}T$ samples from model $i$. We call this the mixture oracle scenario. Nevertheless, in reality, we do not know $\mathbf{K},\mathbf{f}$, and cannot compute $\boldsymbol{\alpha}^{*}$ exactly. Instead, we have to approximate $\boldsymbol{\alpha}^{*}$ by minimizing the sample loss $\hat{L}(\boldsymbol{\alpha};\mathbf{x})$ (5). However, we do not have the samples $\mathbf{x}$ at the beginning in order to compute $\hat{L}(\boldsymbol{\alpha};\mathbf{x})$, so we have to generate some samples first. Yet, to generate these initial samples, we need an estimate of $\boldsymbol{\alpha}^{*}$, or else those samples may have a suboptimal empirical distribution and affect our final loss $L(\hat{P}^{(T)})$, or we will have to discard those initial samples which results in wastage.
This “chicken and egg” problem is naturally solved by an online learning approach via multi-armed bandit. At time $t=1,\ldots,T$, we choose and pull an arm $b^{(t)}\in[m]$ (i.e., generate a sample from model $b^{(t)}$), and obtain a sample $x^{(t)}$ from the distribution $P_{b^{(t)}}$. The choice $b^{(t)}$ can depend on all previous samples $x^{(1)},\ldots,x^{(t-1)}$. Unlike conventional multi-armed bandit where the goal is to maximize the total reward over $T$ rounds, here we minimize the loss $L(\hat{P}^{(T)})$ which involve cross terms $\kappa(x^{(s)},x^{(t)})$ between samples at different rounds. Note that if $\kappa(x,x^{\prime})=0$, then this reduces to the conventional multi-armed bandit setting by taking $f(x)$ to be the negative reward of the sample $x$. In the following subsections, we will propose two new algorithms that are generalizations of the upper confidence bound (UCB) algorithm for multi-armed bandit [18, 19].
@@ 5.1 Mixture Upper Confidence Bound – Continuum-Armed Bandit
We now present the mixture upper confidence bound – continuum-armed bandit (Mixture-UCB-CAB) algorithm. It has a parameter $\beta>1$. It treats the online selection problem as multi-armed bandit with infinitely many arms similar to the continuum-armed bandit settings in [21, 31]. Each arm is a probability vector $\boldsymbol{\alpha}$. By pulling the arm $\boldsymbol{\alpha}$, we generate a sample from a randomly chosen model, where model $i$ is chosen with probability $\alpha_{i}$. Refer to Algorithm 1.
@@
$\boldsymbol{\alpha}^{(t)}:=\mathrm{argmin}_{\boldsymbol{\alpha}}\big{(}\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}\big{)},$ (6)
where the minimization is over probability vectors $\boldsymbol{\alpha}$, and $\boldsymbol{\epsilon}^{(t)}\in\mathbb{R}^{m}$ is defined as
@@
9:     Pull arm $b=b^{(t+1)}$ at time $t+1$ to obtain a new sample $x_{b,n_{b}^{(t)}+1}\sim P_{b}$. Set $n_{b}^{(t+1)}=n_{b}^{(t)}+1$ and $n_{j}^{(t+1)}=n_{j}^{(t)}$ for $j\neq b$33footnotemark: 3
10:end forreturn samples $\mathbf{x}^{(T)}$
Similar to UCB, Mixture-UCB-CAB finds a lower confidence bound $\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}$ of the true loss $L(\boldsymbol{\alpha})$ at each round. To justify the expressions (6), (7), we prove that $\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}$ in (6) lower-bounds $L(\boldsymbol{\alpha})$ with probability at least $1-t^{-\beta}$. The proof is given in Appendix A.1.
Theorem 1.
Fix a probability vector $\boldsymbol{\alpha}$ [^4] Theorem 1 holds for a fixed $\boldsymbol{\alpha}$. A worst-case bound that simultaneously holds for every $\boldsymbol{\alpha}$ is in Lemma 1. Suppose we have samples $x_{i,1},\ldots,x_{i,n_{i}}$ from the distribution $P_{i}$ for $i=1,\ldots,m$, where $n_{i}$ is the number of observed samples from model $i$. For $\delta>0$,
$\mathbb{P}\big{(}\hat{L}(\boldsymbol{\alpha};\mathbf{x})-L(\boldsymbol{\alpha})\geq\boldsymbol{\epsilon}(\delta)^{\top}\boldsymbol{\alpha}\big{)}\leq\delta,\quad\mathbb{P}\big{(}L(\boldsymbol{\alpha})-\hat{L}(\boldsymbol{\alpha};\mathbf{x})\geq\boldsymbol{\epsilon}(\delta)^{\top}\boldsymbol{\alpha}\big{)}\leq\delta,$
where $\boldsymbol{\epsilon}(\delta):=\big{(}\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\big{)}_{i\in[m]}$.
We now prove that Mixture-UCB-CAB gives an expected loss $\mathbb{E}[L(\hat{P}^{(T)})]$ that converges to the optimal loss $\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha})$ by bounding their gap. This means that Mixture-UCB-CAB is a zero-regret strategy by treating $\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha})$ as the average regret per round. [^5] To justify calling $R:=\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha})$ the average regret per round, note that when $\kappa(x,x^{\prime})=0$ and $f(x)=-r(x)$ where $r(x)$ is the reward of the sample $x$, i.e., the loss $L(P)=\mathbb{E}_{X\sim P}[f(X)]$ is linear, $T(\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha}))=T\max_{i\in[m]}\mathbb{E}_{X\sim P_{i}}[r(X)]-\mathbb{E}[\sum_{t=1}^{T}r(x^{(t)})]$ indeed reduces to the conventional notion of regret. So $R$ can be regarded as the quadratic generalization of regret. The proof is given in Appendix A.2.
Theorem 2.
@@
$\mathbb{E}\big{[}L(\hat{P}^{(T)})\big{]}-\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha})$ $\leq 4\Delta_{L}\sqrt{\frac{\beta m\log T}{T}}.$
When $\kappa(x,x^{\prime})=0$, Mixture-UCB-CAB reduces to the conventional UCB, and Theorem 2 coincides with the $O(\sqrt{(m\log T)/T})$ distribution-free bound on the regret per round of conventional UCB [20]. Since there is a $\Omega(\sqrt{m/T})$ minimax lower bound on the regret per round even for conventional multi-armed bandit without the quadratic kernel term [20, Theorem 3.4], Theorem 2 is tight up to a logarithmic factor.
The main difference between Mixture-UCB-CAB and conventional UCB is that we choose a mixture of arms in (6) given by the probability vector $\boldsymbol{\alpha}$, instead of a single arm. A more straightforward application of UCB would be to simply find the single arm that minimizes the lower bound in (6), i.e., we restrict $\boldsymbol{\alpha}=\mathbf{e}_{i}$ for some $i\in[m]$, where $\mathbf{e}_{i}$ is the $i$-th basis vector, and minimize (6) over $i$ instead. We call this Vanilla-UCB. Vanilla-UCB fails to take into account the possibility that a mixture may give a smaller loss than every single arm. In the long run, Vanilla-UCB converges to pulling the best single arm instead of the optimal mixture. Vanilla-UCB will be used as a baseline to be compared with Mixture-UCB-CAB, and another new algorithm presented in the next section.
Mixture-UCB-CAB can be extended to the Sparse-Mixture-UCB-CAB algorithm which eventually select only a small number of models. This can be useful if there is a subscription cost for each model. Refer to Appendix B for discussions.
55footnotetext: We may also consider the scenario where each pull gives a batch of $l$ samples instead of only one sample. In this case, we will have $x_{b,n_{b}^{(t-1)}+1},\ldots,x_{b,n_{b}^{(t-1)}+l}\sim P_{b}$ and $n_{b}^{(t)}=n_{b}^{(t-1)}+l$.
@@ 5.2 Mixture Upper Confidence Bound – Online Gradient Descent
We present an alternative to Mixture-UCB-CAB, called the mixture upper confidence bound – online gradient descent (Mixture-UCB-OGD) algorithm, inspired by the online gradient descent algorithm [32]. It also has a parameter $\beta>1$. Refer to Algorithm 2.
@@
$\mathbf{h}^{(t)}$ $:=\nabla_{\boldsymbol{\alpha}}\left(\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}\right)\Big{|}_{\boldsymbol{\alpha}=\mathbf{n}^{(t)}/t}\;=\;\frac{2}{t}\hat{\mathbf{K}}(\mathbf{x}^{(t)})\mathbf{n}^{(t)}+\hat{\mathbf{f}}(\mathbf{x}^{(t)})-\boldsymbol{\epsilon}^{(t)},$ (8)
where $\mathbf{n}^{(t)}:=(n_{i}^{(t)})_{i\in[m]}\in\mathbb{R}^{m}$, and $\boldsymbol{\epsilon}^{(t)}$ is defined as in Mixture-UCB-CAB
@@
9:end forreturn samples $\mathbf{x}^{(T)}$
@@
Both Mixture-UCB-CAB and Mixture-UCB-OGD attempt to make the “proportion vector” $\mathbf{n}^{(t)}/t$ (note that $n_{i}^{(t)}/t$ is the proportion of samples from model $i$) approach the optimal mixture $\boldsymbol{\alpha}^{*}$ that minimizes $L(\boldsymbol{\alpha})$, but they do so in different manners. Mixture-UCB-CAB first computes the estimate $\boldsymbol{\alpha}^{(t)}$ after time $t$, then approaches $\boldsymbol{\alpha}^{(t)}$ by pulling an arm randomly chosen from the distribution $\boldsymbol{\alpha}^{(t)}$. Mixture-UCB-OGD estimates the gradient $\mathbf{h}^{(t)}$ of the loss function at the current proportion vector $\mathbf{n}^{(t)}/t$, and pulls an arm that results in the steepest descent of the loss.

### v2

@@ 4 Optimal Mixtures of Generative Models
Suppose we are given $m$ generative models, where model $i$ generates samples from the distribution $P_{i}$. If our goal is merely to find the model that minimize the loss (4), we should select $\mathrm{argmin}_{i}L(P_{i})$. Nevertheless, for diversity metrics such as RKE, it is possible that a mixture of the models will give a better diversity. Assume that the mixture weight of model $i$ is $\alpha_{i}\in[0,1]$, where $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{m})$ is a probability vector. The loss of the mixture distribution $\sum_{i=1}^{m}\alpha_{i}P_{i}$ can then be expressed as
$L(\bm{\alpha})$ $:=L\Big{(}\sum_{i=1}^{m}\alpha_{i}P_{i}\Big{)}=\bm{\alpha}^{\top}\mathbf{K}\bm{\alpha}+\mathbf{f}^{\top}\bm{\alpha},$
@@
Given $\mathbf{K},\mathbf{f}$, the probability vector $\bm{\alpha}$ minimizing $L(\bm{\alpha})$ can be found via a convex quadratic program.
In practice, we do not know the precise $\mathbf{K},\mathbf{f}$, and have to estimate them using samples. Suppose we have the samples $x_{i,1},\ldots,x_{i,n_{i}}$ from the distribution $P_{i}$ for $i=1,\ldots,m$, where $n_{i}$ is the number of observed samples from model $i$. Write $\mathbf{x}:=(x_{i,a})_{i\in[m],\,a\in[n_{i}]}$. We approximate the true mixture distribution $\sum_{i=1}^{m}\alpha_{i}P_{i}$ by the empirical mixture distribution $\sum_{i=1}^{m}\frac{\alpha_{i}}{n_{i}}\sum_{a=1}^{n_{i}}\delta_{x_{i,a}}$, where we assign a weight $\alpha_{i}/n_{i}$ to samples $x_{i,a}$ from model $i$, and $\delta_{x_{i,a}}$ denotes the degenerate distribution at $x_{i,a}$. We then approximate $L(\bm{\alpha})$ by the sample loss
$\hat{L}(\bm{\alpha};\mathbf{x})$ $:=L\Big{(}\sum_{i=1}^{m}\frac{\alpha_{i}}{n_{i}}\sum_{a=1}^{n_{i}}\delta_{x_{i,a}}\Big{)}=\bm{\alpha}^{\top}\hat{\mathbf{K}}(\mathbf{x})\bm{\alpha}+\hat{\mathbf{f}}(\mathbf{x})^{\top}\bm{\alpha},$ (5)
@@
The minimization of $\hat{L}(\bm{\alpha};\mathbf{x})$ over probability vectors $\bm{\alpha}$ is still a convex quadratic program.
@@ 5 Online Selection of Optimal Mixtures – Mixture Multi-Armed Bandit
If we are told by an oracle the optimal mixture $\bm{\alpha}^{*}$ that minimizes the loss $L(\bm{\alpha})$, then we should generate samples according to this mixture distribution, giving $\approx\alpha^{*}_{i}T$ samples from model $i$. We call this the mixture oracle scenario. Nevertheless, in reality, we do not know $\mathbf{K},\mathbf{f}$, and cannot compute $\bm{\alpha}^{*}$ exactly. Instead, we have to approximate $\bm{\alpha}^{*}$ by minimizing the sample loss $\hat{L}(\bm{\alpha};\mathbf{x})$ (5). However, we do not have the samples $\mathbf{x}$ at the beginning in order to compute $\hat{L}(\bm{\alpha};\mathbf{x})$, so we have to generate some samples first. Yet, to generate these initial samples, we need an estimate of $\bm{\alpha}^{*}$, or else those samples may have a suboptimal empirical distribution and affect our final loss $L(\hat{P}^{(T)})$, or we will have to discard those initial samples which results in wastage.
This “chicken and egg” problem is naturally solved by an online learning approach via multi-armed bandit. At time $t=1,\ldots,T$, we choose and pull an arm $b^{(t)}\in[m]$ (i.e., generate a sample from model $b^{(t)}$), and obtain a sample $x^{(t)}$ from the distribution $P_{b^{(t)}}$. The choice $b^{(t)}$ can depend on all previous samples $x^{(1)},\ldots,x^{(t-1)}$. Unlike conventional multi-armed bandit where the goal is to maximize the total reward over $T$ rounds, here we minimize the loss $L(\hat{P}^{(T)})$ which involve cross terms $\kappa(x^{(s)},x^{(t)})$ between samples at different rounds. Note that if $\kappa(x,x^{\prime})=0$, then this reduces to the conventional multi-armed bandit setting by taking $f(x)$ to be the negative reward of the sample $x$. In the following subsections, we will propose two new algorithms that are generalizations of the upper confidence bound (UCB) algorithm for multi-armed bandit (Agrawal, 1995a; Auer, 2002).
@@ 5.1 Mixture Upper Confidence Bound – Continuum-Armed Bandit
We now present the mixture upper confidence bound – continuum-armed bandit (Mixture-UCB-CAB) algorithm. It has a parameter $\beta>1$. It treats the online selection problem as multi-armed bandit with infinitely many arms similar to the continuum-armed bandit settings in (Agrawal, 1995b; Lu et al., 2019). Each arm is a probability vector $\bm{\alpha}$. By pulling the arm $\bm{\alpha}$, we generate a sample from a randomly chosen model, where model $i$ is chosen with probability $\alpha_{i}$. Refer to Algorithm 1.
@@
$\bm{\alpha}^{(t)}:=\mathrm{argmin}_{\bm{\alpha}}\big{(}\hat{L}(\bm{\alpha};\mathbf{x}^{(t)})-(\bm{\epsilon}^{(t)})^{\top}\bm{\alpha}\big{)},$ (6)
where the minimization is over probability vectors $\bm{\alpha}$, and $\bm{\epsilon}^{(t)}\in\mathbb{R}^{m}$ is defined as
@@
9:     Pull arm $b=b^{(t+1)}$ at time $t+1$ to obtain a new sample $x_{b,n_{b}^{(t)}+1}\sim P_{b}$. Set $n_{b}^{(t+1)}=n_{b}^{(t)}+1$ and $n_{j}^{(t+1)}=n_{j}^{(t)}$ for $j\neq b$55footnotemark: 5
10:  end for
11:  return  samples $\mathbf{x}^{(T)}$
Similar to UCB, Mixture-UCB-CAB finds a lower confidence bound $\hat{L}(\bm{\alpha};\mathbf{x}^{(t)})-(\bm{\epsilon}^{(t)})^{\top}\bm{\alpha}$ of the true loss $L(\bm{\alpha})$ at each round. To justify the expressions (6), (7), we prove that $\hat{L}(\bm{\alpha};\mathbf{x}^{(t)})-(\bm{\epsilon}^{(t)})^{\top}\bm{\alpha}$ in (6) lower-bounds $L(\bm{\alpha})$ with probability at least $1-t^{-\beta}$. The proof is given in Appendix 8.1.
Theorem 1
Fix a probability vector $\bm{\alpha}$ [^6] Theorem 1 holds for a fixed $\bm{\alpha}$. A worst-case bound that simultaneously holds for every $\bm{\alpha}$ is in Lemma 1. Suppose we have samples $x_{i,1},\ldots,x_{i,n_{i}}$ from the distribution $P_{i}$ for $i=1,\ldots,m$, where $n_{i}$ is the number of observed samples from model $i$. For $\delta>0$,
$\mathbb{P}\big{(}\hat{L}(\bm{\alpha};\mathbf{x})-L(\bm{\alpha})\geq\bm{\epsilon}(\delta)^{\top}\bm{\alpha}\big{)}\leq\delta,\quad\mathbb{P}\big{(}L(\bm{\alpha})-\hat{L}(\bm{\alpha};\mathbf{x})\geq\bm{\epsilon}(\delta)^{\top}\bm{\alpha}\big{)}\leq\delta,$
where $\bm{\epsilon}(\delta):=\big{(}\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\big{)}_{i\in[m]}$.
We now prove that Mixture-UCB-CAB gives an expected loss $\mathbb{E}[L(\hat{P}^{(T)})]$ that converges to the optimal loss $\min_{\bm{\alpha}}L(\bm{\alpha})$ by bounding their gap. This means that Mixture-UCB-CAB is a zero-regret strategy by treating $\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\bm{\alpha}}L(\bm{\alpha})$ as the average regret per round. [^7] To justify calling $R:=\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\bm{\alpha}}L(\bm{\alpha})$ the average regret per round, note that when $\kappa(x,x^{\prime})=0$ and $f(x)=-r(x)$ where $r(x)$ is the reward of the sample $x$, i.e., the loss $L(P)=\mathbb{E}_{X\sim P}[f(X)]$ is linear, $T(\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\bm{\alpha}}L(\bm{\alpha}))=T\max_{i\in[m]}\mathbb{E}_{X\sim P_{i}}[r(X)]-\mathbb{E}[\sum_{t=1}^{T}r(x^{(t)})]$ indeed reduces to the conventional notion of regret. So $R$ can be regarded as the quadratic generalization of regret. The proof is given in Appendix 8.2.
Theorem 2
@@
$\mathbb{E}\big{[}L(\hat{P}^{(T)})\big{]}-\min_{\bm{\alpha}}L(\bm{\alpha})$ $\leq 4\Delta_{L}\sqrt{\frac{\beta m\log T}{T}}.$
When $\kappa(x,x^{\prime})=0$, Mixture-UCB-CAB reduces to the conventional UCB, and Theorem 2 coincides with the $O(\sqrt{(m\log T)/T})$ distribution-free bound on the regret per round of conventional UCB (Bubeck et al., 2012). Since there is a $\Omega(\sqrt{m/T})$ minimax lower bound on the regret per round even for conventional multi-armed bandit without the quadratic kernel term (Bubeck et al., 2012, Theorem 3.4), Theorem 2 is tight up to a logarithmic factor.
The main difference between Mixture-UCB-CAB and conventional UCB is that we choose a mixture of arms in (6) given by the probability vector $\bm{\alpha}$, instead of a single arm. A more straightforward application of UCB would be to simply find the single arm that minimizes the lower bound in (6), i.e., we restrict $\bm{\alpha}=\mathbf{e}_{i}$ for some $i\in[m]$, where $\mathbf{e}_{i}$ is the $i$-th basis vector, and minimize (6) over $i$ instead. We call this Vanilla-UCB. Vanilla-UCB fails to take into account the possibility that a mixture may give a smaller loss than every single arm. In the long run, Vanilla-UCB converges to pulling the best single arm instead of the optimal mixture. Vanilla-UCB will be used as a baseline to be compared with Mixture-UCB-CAB, and another new algorithm presented in the next section.
Mixture-UCB-CAB can be extended to the Sparse-Mixture-UCB-CAB algorithm which eventually select only a small number of models. This can be useful if there is a subscription cost for each model. Refer to Appendix 8.3 for discussions.
77footnotetext: We may also consider the scenario where each pull gives a batch of $l$ samples instead of only one sample. In this case, we will have $x_{b,n_{b}^{(t-1)}+1},\ldots,x_{b,n_{b}^{(t-1)}+l}\sim P_{b}$ and $n_{b}^{(t)}=n_{b}^{(t-1)}+l$.
@@ 5.2 Mixture Upper Confidence Bound – Online Gradient Descent
We present an alternative to Mixture-UCB-CAB, called the mixture upper confidence bound – online gradient descent (Mixture-UCB-OGD) algorithm, inspired by the online gradient descent algorithm (Shalev-Shwartz et al., 2012). It also has a parameter $\beta>1$. Refer to Algorithm 2.
@@
$\mathbf{h}^{(t)}$ $:=\nabla_{\bm{\alpha}}\left(\hat{L}(\bm{\alpha};\mathbf{x}^{(t)})-(\bm{\epsilon}^{(t)})^{\top}\bm{\alpha}\right)\Big{|}_{\bm{\alpha}=\mathbf{n}^{(t)}/t}\;=\;\frac{2}{t}\hat{\mathbf{K}}(\mathbf{x}^{(t)})\mathbf{n}^{(t)}+\hat{\mathbf{f}}(\mathbf{x}^{(t)})-\bm{\epsilon}^{(t)},$ (8)
where $\mathbf{n}^{(t)}:=(n_{i}^{(t)})_{i\in[m]}\in\mathbb{R}^{m}$, and $\bm{\epsilon}^{(t)}$ is defined as in Mixture-UCB-CAB
@@
9:  end for
10:  return  samples $\mathbf{x}^{(T)}$
@@
Both Mixture-UCB-CAB and Mixture-UCB-OGD attempt to make the “proportion vector” $\mathbf{n}^{(t)}/t$ (note that $n_{i}^{(t)}/t$ is the proportion of samples from model $i$) approach the optimal mixture $\bm{\alpha}^{*}$ that minimizes $L(\bm{\alpha})$, but they do so in different manners. Mixture-UCB-CAB first computes the estimate $\bm{\alpha}^{(t)}$ after time $t$, then approaches $\bm{\alpha}^{(t)}$ by pulling an arm randomly chosen from the distribution $\bm{\alpha}^{(t)}$. Mixture-UCB-OGD estimates the gradient $\mathbf{h}^{(t)}$ of the loss function at the current proportion vector $\mathbf{n}^{(t)}/t$, and pulls an arm that results in the steepest descent of the loss.